OpenStudy (anonymous):

log(3)(3x+2)=4 solve for x

OpenStudy (dumbcow):

remember move the base over and make the 4 the exponent 3x +2 = 3^4 3x+2 = 81

OpenStudy (anonymous):

26.333?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

if you look carefully at the problem, you can see that x is roughly 1 million trillion thousand pi